The sum of 3 consecutive integers is 13 more than twice the smallest integer. Find the integer.
Answer by ankor@dixie-net.com(12684)

The sum of 3 consecutive integers is 13 more than twice the smallest integer. Find the integer.
:
x + (x+1) + (x+2) = 2x + 13
:
3x + 3 = 2x + 13
:
3x - 2x = 13 - 3
:
x = 10 is the smallest integer:
:
:
Check: 10 + 11 + 12 = 2(10) + 13